# Sensitivity Analysis of Image Classification Models using Generalized Polynomial Chaos

### v1

@@ 2.2 Generalized polynomial chaos
[…]
@@
[…]
@@ 3 GPC-SA FOR IMAGE CLASSIFICATION MODELS
This approach allows us to isolate and analyze the influence of the perturbations $\boldsymbol{\xi}$ on the model’s predictions, facilitated by the computation of Sobol indices to quantify the influence of $\boldsymbol{\xi}$ on the model’s output. However, the range of polynomials is (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), while the range of a discrete probability distribution is $[0,1]$. To account for this, we apply the logit function in order to map the probability to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). This can be written as
@@ 4.1 TIG welding seam classification
Figure 2: Figure (A) presents a surface plot of the GPC approximation, illustrating the relationship between brightness and rotation for the prediction probability, with tilt held constant at 00. Figure (B) displays the approximation of brightness while both rotation and tilt are fixed at 00.
@@
For this purpose, we draw $n=1000$ samples using Latin hypercube sampling (LHS) from a Beta distribution with specified limits for each transformation parameter. The limits are $[0,2]$ for brightness (representing the scaling factor of brightness change, where $1$ represents the original brightness), $[-30,30]$ for rotation (the angle of rotation in degrees), and $[-20,20]$ for tilt (the angle of tilt in degrees).
@@
Table 1 presents the resulting Sobol indices. These indices indicate that tilt has the highest influence on the model, with a value of $0.37$. In contrast, brightness and rotation exert the lowest influence on the model, each with a value of $0.03$. The combined relative influence of brightness, rotation, and tilt is 0.0610.0610.0610.061111For those interested in reproducing the results or exploring the SA approach further, the code including the fine-tuned ResNet18 model is available on GitHub at https://github.com/lpossner/GPC_Image_Classification..
The surrogate models illustrating the GPC approximation for rotation and brightness are depicted in Figure 2. Figure 2 (A) presents a surface plot of the surrogate model with respect to rotation and brightness, where tilt is fixed at a value of 00. Figure 2 (B) shows a plot of the surrogate model with respect to brightness, where rotation and tilt are fixed at a value of 00.
@@ 4.2 Vehicle emblem classification
To conduct our experiment, we set up an apparatus comprising a Keyence camera, an adjustable light source, a luminance meter, and an angle scale. We then adjusted the rotation and brightness for $n=100$ samples using LHS drawn from a Beta distribution. The limits for the rotation transformation (the angle of rotation in degrees) are set between $[-180,180]$, and for the brightness transformation (representing brightness in Lux), the range is $[0,7000]$. With this configuration, we adjust the rotation angle and brightness and shot the images (cf. Figure 3 (B)) to compute the output class probabilities using Equation (7). To approximate the surrogate model, we employ GPC with Jacobian polynomials and subsequently compute the Sobol indices. The relative error obtained from this method is $0.23$, given a maximum total polynomial order of $[6,5]$.

### v2

@@ 2.2 Generalized polynomial chaos
$\int\psi_{\alpha_{i}}(\xi_{i})\psi_{\beta_{i}}(\xi_{i})p_{\Xi_{i}}(\xi_{i})\,d\xi_{i}=\lVert\psi_{\alpha_{i}}(\xi_{i})\rVert^{2}\delta_{\alpha_{i},\beta_{i}},$ (2)
@@
$c_{\boldsymbol{\alpha}}=\frac{1}{\prod_{i}\lVert\psi_{i}(\xi_{i})\rVert^{2}}\int f(\boldsymbol{\xi})\Psi(\boldsymbol{\xi})p_{\Xi}(\boldsymbol{\xi})\,d\boldsymbol{\xi}.$ (3)
@@ 3 GPC-SA FOR IMAGE CLASSIFICATION MODELS
This approach allows us to isolate and analyze the influence of the perturbations $\boldsymbol{\xi}$ on the model’s predictions, facilitated by the computation of Sobol indices to quantify the influence of $\boldsymbol{\xi}$ on the model’s output. However, the range of polynomials is $(-\infty,+\infty)$, while the range of a discrete probability distribution is $[0,1]$. To account for this, we apply the logit function in order to map the probability to $(-\infty,+\infty)$. This can be written as
@@ 4.1 TIG welding seam classification
Figure 2: Figure (A) presents a surface plot of the GPC approximation, illustrating the relationship between brightness and rotation for the prediction probability, with tilt held constant at 0. Figure (B) displays the approximation of brightness while both rotation and tilt are fixed at 0.
@@
For this purpose, we draw $n=1000$ samples using Latin hypercube sampling (LHS) from a Beta distribution with specified limits for each transformation parameter. The limits are $[0,2]$ for brightness (representing the scaling factor of brightness change, where 11 represents the original brightness), [30,30][-30,30] for rotation (the angle of rotation in degrees), and [20,20][-20,20] for tilt (the angle of tilt in degrees).
@@
Table 1 presents the resulting Sobol indices. These indices indicate that tilt has the highest influence on the model, with a value of $0.37$. In contrast, brightness and rotation exert the lowest influence on the model, each with a value of $0.03$. The combined relative influence of brightness, rotation, and tilt is 0.0610.061111For those interested in reproducing the results or exploring the SA approach further, the code including the fine-tuned ResNet18 model is available on GitHub at https://github.com/lpossner/GPC_Image_Classification..
The surrogate models illustrating the GPC approximation for rotation and brightness are depicted in Figure 2. Figure 2 (A) presents a surface plot of the surrogate model with respect to rotation and brightness, where tilt is fixed at a value of 0. Figure 2 (B) shows a plot of the surrogate model with respect to brightness, where rotation and tilt are fixed at a value of 0.
@@ 4.2 Vehicle emblem classification
To conduct our experiment, we set up an apparatus comprising a Keyence camera, an adjustable light source, a luminance meter, and an angle scale. We then adjusted the rotation and brightness for $n=100$ samples using LHS drawn from a Beta distribution. The limits for the rotation transformation (the angle of rotation in degrees) are set between [180,180][-180,180], and for the brightness transformation (representing brightness in Lux), the range is $[0,7000]$. With this configuration, we adjust the rotation angle and brightness and shot the images (cf. Figure 3 (B)) to compute the output class probabilities using Equation (7). To approximate the surrogate model, we employ GPC with Jacobian polynomials and subsequently compute the Sobol indices. The relative error obtained from this method is $0.23$, given a maximum total polynomial order of $[6,5]$.